the length of a rectangle is 3 times the width. the difference between the lenght and the width is 8 feet determine the rectangle?
Answer by checkley77(12844)

L=3W
L-W=8
3W-W=8
2W=8
W=8/2
W=4 ANS. FOR THE WIDTH.
L=3*4=12 ANS. FOR THE LENGTH.